OpenStudy (anonymous):

solve for x: log64=2logx

OpenStudy (anonymous):

log64 = logx^2 64 = x^2 x = 8 or -8 (Your instructor may want you to rule out the -8)

OpenStudy (anonymous):

thx ur rlly hepfull how do i become ur fan

OpenStudy (anonymous):

Haha I don't know, I just joined

OpenStudy (anonymous):

same here